OpenStudy (anonymous):

Find the area under the curve y=5xe^{-x} for x ≥ 1

OpenStudy (anonymous):

Have you learned integration by parts?

OpenStudy (science0229):

Actually, do you know improper integral?

OpenStudy (science0229):

Never mind my last comment. If you know limit, you should be fine while I take you step by step.

OpenStudy (science0229):

First, let's find a general formula for the area under curve in the interval [1,t]

OpenStudy (science0229):

\[\int\limits_{1}^{t}5xe^{-x}\]

OpenStudy (science0229):

Now, as @TanteAnna asked, do you know integration by parts?

OpenStudy (anonymous):

\[\int\limits_1^{\infty } 5 x e^{-x} \, dx=\frac{10}{e} \]

OpenStudy (anonymous):

Thanks guys